OpenStudy (anonymous):

A right triangle has legs of 4 and 5. What is the hypotenuse? A right triangle has a hypotenuse of 13 and a leg of 8. What is the other leg? -I dont understand this :(

OpenStudy (aripotta):

Pythagorean theorem: a^2 + b^2 = c^2. a abd b represents the two legs, and c represents the hypotenuse

OpenStudy (aripotta):

a and b*

OpenStudy (aripotta):

4^2 + 5^2 = c^2. first find out what 4^2 and 5^2 is

OpenStudy (kymber):

You need to use the Pythagorean Theorem, which is a² + b² = c² c is the hypotenuse, the longest side, and a + b are the legs. You plug in what you know, then solve For the first one: 4² + 5² = c² The second one: a² + 8² = 13² Now just solve it \(\huge\ddot\smile\)

OpenStudy (anonymous):

for the first one i got 41?

OpenStudy (aripotta):

yea, that when you add 4^2 and 5^2 but now you have to find the square root of 41 to get your hypotenuse

OpenStudy (aripotta):

that's*

OpenStudy (anonymous):

6.4

OpenStudy (aripotta):

yup

OpenStudy (aripotta):

do you know how to do the second one?

OpenStudy (anonymous):

nope

OpenStudy (aripotta):

ok, so the theorem is a^2 + b^2 = c^2. we know what one of the legs is, a, and what the hypotenuse is, c. so after we put in the numbers, our equation will look like this: 8^2 + b^2 = 13^2

OpenStudy (aripotta):

so 64 + b^2 = 169

OpenStudy (aripotta):

how do you solve for b?

OpenStudy (anonymous):

i was always told to subtract

OpenStudy (aripotta):

yea, so subtract 64 from 169

OpenStudy (aripotta):

then find the square root

OpenStudy (anonymous):

10.2

OpenStudy (aripotta):

yep, so there's your answer. do you understand what you did to get it?